OpenStudy (imnotgoodatmath):

Simplify (12)^-3 . (12)^-5

OpenStudy (anonymous):

use this x^3*x^n=x^(m+n)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 12^{-3} \times 12^{-5} }\) like this ?

OpenStudy (anonymous):

oh..its m..not 3..!!!

OpenStudy (imnotgoodatmath):

the answers are 11 . -2 1/11.11 1/11 . 2 . 2 -11 . 2. 2

OpenStudy (imnotgoodatmath):

yes @SolomonZelman

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ 12^{-3} \times 12^{-5} =12^{\LARGE-3+(-5)} =12^{-3-5}=12^{-8}=\frac{1}{12^{\LARGE8} }=\frac{1}{429981696} }\)

OpenStudy (solomonzelman):

You can call this number zero (unofficially) if you know what I mean:)

OpenStudy (imnotgoodatmath):

dang lol

OpenStudy (solomonzelman):

;)

OpenStudy (imnotgoodatmath):

so what you think it is lol?

OpenStudy (solomonzelman):

your choices are unclear.

OpenStudy (imnotgoodatmath):

dang you school

OpenStudy (imnotgoodatmath):

Simplify twelve to the negative third times twelve to the negative fifth. 12 - 2 12 - 8

OpenStudy (imnotgoodatmath):

12^8 12^15

OpenStudy (solomonzelman):

you mean 1/12^8 and 1/12^15 ?

OpenStudy (imnotgoodatmath):

i dont get it the ones i put were the answer choices

OpenStudy (solomonzelman):

it is 12^-8

OpenStudy (imnotgoodatmath):

ok thx just asking but what about the top one

OpenStudy (imnotgoodatmath):

@SolomonZelman